OpenStudy (anonymous):

f(x)=2(x-2)^2-4 Find the Vertex

OpenStudy (turingtest):

the general form of a parabola with vertex\[(h,k)\]is\[y-k=a(x-h)^2\]put your equation in that form and find h and k, then you have your answer.

OpenStudy (turingtest):

btw here y=f(x)